OpenStudy (anonymous):

Given that tan a = -sqrt5 / 2 , where a is an obtuse angle, find the value of sin a/2 without using calculator

OpenStudy (anonymous):

Well \[\Large tan \theta=\frac{sin\theta}{cos\theta}\] So you have \[\Large \frac{-\sqrt{5}}{2}\] Well you have the same situation here where you can find the sin

OpenStudy (anonymous):

Sin=numerator Cos=denominator

jhonyy9 (jhonyy9):

do you have any idea to calculi it without calculator ?

OpenStudy (anonymous):

He didn't see this prob yet but @jhonyy9 was that a worng approach to guiding ihm to hte answer?

OpenStudy (anonymous):

him to the *

jhonyy9 (jhonyy9):

ok ,you have any idea to calculi it ,?so than how ?

OpenStudy (anonymous):

Are you talking to me?

jhonyy9 (jhonyy9):

yes sure

OpenStudy (anonymous):

Unless he made a typo with writng sin a/2.....Then the sin is -sqrt5 cause the cos which is 2, you divide it to get the tan

jhonyy9 (jhonyy9):

in what quadrant will has cosinus value of 2 ,do you know it ?

OpenStudy (anonymous):

Well this has to be in the 4th quadrant since sin is negative and cos is positive and in order for cos to be positive, both sin/cos have to be pos or sin is negative alone

OpenStudy (anonymous):